Helium gas in a balloon occupies 2.5 L at 300.0 K. The balloon is dipped into liquid nitrogen that is at a temperature of 80.0 k. What will the volume of the helium in the balloon at the lower temperature be?

I got a weird answer like .66667 and i used charles law
Answer #1
assuming pressure would be the same (it is not if one is submerged in a liquid).

V1/T1=V2/T2

V2=2.5*80/(300)=2.5*4/15 which is about in my head 1/4*2.5 which is your answer
answered by: duppala
Answer #2
thanku!
answered by: lele